The average price of a TI 89 titanium calculator with the term to be 152.33 from a random sample of nine calculators being sold online assume that the population is normally distributed in the population third vision is $20 Quesly in 95 all TI 89 being sold online


The average price of a TI 89 titanium calculator with the term to be 152.33 from a random sample of nine calculators being sold online assume that the population is normally distributed in the population stabdard deviation $28 calculate a 95% cobfidence interval for the mean price of ll ti-89 sold online
Answer #1

sample mean, xbar = 152.33
sample standard deviation, σ = 28
sample size, n = 9


Given CI level is 95%, hence α = 1 - 0.95 = 0.05
α/2 = 0.05/2 = 0.025, Zc = Z(α/2) = 1.96


ME = zc * σ/sqrt(n)
ME = 1.96 * 28/sqrt(9)
ME = 18.29

CI = (xbar - Zc * s/sqrt(n) , xbar + Zc * s/sqrt(n))
CI = (152.33 - 1.96 * 28/sqrt(9) , 152.33 + 1.96 * 28/sqrt(9))
CI = (134.04 , 170.62)